# Enhancing Robustness of Vision-Language Models through Orthogonality Learning and Cross-Regularization

### v1

@@ Abstract
Efficient finetuning of vision-language models (VLMs) like CLIP for specific downstream tasks is gaining significant attention. Previous works primarily focus on prompt learning to adapt the CLIP into a variety of downstream tasks, however, suffering from task overfitting when finetuned on a small data set. In this paper, we introduce an orthogonal finetuning method for efficiently updating pretrained weights which enhances robustness and generalization, while a cross-regularization strategy is further exploited to maintain the stability in terms of zero-shot generalization of VLMs, dubbed OrthCR. Specifically, trainable orthogonal matrices are injected seamlessly into the transformer architecture and enforced with orthogonality constraint using Cayley parameterization, benefiting from the norm-preserving property and thus leading to stable and faster convergence. To alleviate deviation from orthogonal constraint during training, a cross-regularization strategy is further employed with initial pretrained weights within a bypass manner. In addition, to enrich the sample diversity for downstream tasks, we first explore Cutout data augmentation to boost the efficient finetuning and comprehend how our approach improves the specific downstream performance and maintains the generalizability in the perspective of Orthogonality Learning. Beyond existing prompt learning techniques, we conduct extensive experiments to demonstrate that our method explicitly steers pretrained weight space to represent the task-specific knowledge and presents competitive generalizability under base-to-base/base-to-new, cross-dataset transfer and domain generalization evaluations.
@@ 1 Introduction
Large-scale pretrained vision-language models (VLMs) have been emerging as prevalent cornerstones in a wide spectrum of downstream vision and vision-language tasks, including few-shot image recognition [87; 88; 85; 21; 37; 89; 68; 56; 11; 74], object-detection [20; 24; 3; 82] and segmentation [17; 6; 65; 76]. Leading models like CLIP [64] and ALIGN [35] demonstrate remarkable generalizability by aligning image-text pairs from large web corpora using contrastive loss, thereby encoding open-vocabulary concepts within a joint vision-language embedding space. Despite the effectiveness of these VLMs in zero-shot recognition, finetuning them for specific downstream tasks while preserving their strong zero-shot capabilities remains a significant challenge. Designing manual text prompts for different tasks requires substantial human effort and expert knowledge, which is often infeasible for achieving optimal performance in data-efficient settings [8].
@@
Recently, prompt learning [88; 87] serves as an exceptional paradigm to achieve this objective, however, tending to prioritize task-specific knowledge and resulting in task overfitting issues [59; 38], where the fine-tuned model struggles to generalize well to new/unseen tasks under data-efficient settings. To address this dilemma, alternative approaches must be explored. Drawing inspiration from empirical observations that hyperspherical similarity effectively encodes semantic information [9; 52; 53] and that hyperspherical energy [50] can characterize the pairwise relational structure among neurons, we hypothesize that well-pretrained models like CLIP should maintain consistent levels of hyperspherical energy even after finetuning. An intuitive approach is to use a suitable regularizer to preserve hyperspherical energy levels during the finetuning phase. However, ensuring that the difference in hyperspherical energy is minimized remains a challenge. Inspired by recent orthogonal transformation methods [63; 51], we propose that the pretrained pairwise hyperspherical energy can be preserved by leveraging orthogonal transformation for all neurons with the same operation. This approach utilizes the invariance property of orthogonal transformation, meaning norm-preserving during finetuning, to maintain consistent hyperspherical energy levels.
@@
Motivated by the preservation of hyperspherical energy through orthogonal transformation, we introduce Orthogonality Learning to adapt pretrained VLMs (e.g., CLIP) to specific downstream tasks (e.g., few-shot image recognition) without altering their hyperspherical energy, thanks to the norm-preserving property during finetuning. This approach differs from common methods that rely on prompt learning. Furthermore, previous works [47; 50; 51] have shown that small hyperspherical energy leads to better generalization, and orthogonal transformation is a suitable and flexible solution for achieving this, especially in classification task. Our main idea is to apply the same orthogonal transformation to neurons so that pairwise angles are maintained within the hypersphere of CLIP. Although prevalent adaptation methods for pretrained weights, such as LoRA [32], achieve finetuning by adding small component matrices, they still suffer from low training convergence and generalizability degradation.
Figure 1: The pipeline comparison for tuning or adapting VLMs into downstream tasks. Our contribution is to introduce a new finetuning pipeline by orthogonal tuning.
@@
In this paper, we propose a novel and efficient finetuning method using Orthogonality Learning, motivated by the preservation of hyperspherical energy through orthogonal transformation, shown different paradigm with exisitng works in Fig 1. To mitigate deviation from orthogonal constraint during training, we introduce a Cross-Regularization strategy using the initial pretrained weights as an anchor point, thus dubbed OrthCR. Our method keeps the pretrained weights frozen while applying orthogonal finetuning and regularization simultaneously. In the dual-branch transformer architecture of the CLIP model, we inject trainable orthogonal matrices and enforce orthogonal constraints using Cayley parameterization [28; 42]. This ensures each injected layer matrix is orthogonal with a determinant of 1. We investigate orthogonal finetuning in both image and text encoder of CLIP to demonstrate training efficiency and generalizability preservation of our method, distinguishing it from prompt tuning and low-rank matrix decomposition methods. The norm-preserving property of orthogonal transformations helps maintain hyperspherical energy levels, benefiting of stable convergence, robustness, and generalization. This enables seamless integration of task-specific knowledge into pretrained VLMs, allowing the trainable matrices to be merged with frozen weights during deployment without adding inference latency. To prevent significant deviations from the pretrained model, we employ a Cross-Regularization strategy that guides the model to stay close to the anchor point, supported by the pretrained model within a bypass manner. Instead of explicitly minimizing the distance between finetuned and pretrained weights, we use a logit distillation loss to regularize the logits from the finetuned text encoder while detaching updates from image encoder, and vice versa. This simple yet effective approach sustains orthogonal finetuning with initial anchor regularization, avoiding deviations from the zero-shot generalizability manifold severely. Besides, we utilize Cutout data augmentation to enrich the data diversity, enhancing the task-specific knowledge of finetuned model (e.g., few-shot image recognition) under data-efficient setting. Unlike previous works [63; 51], we focus on adapting VLMs to high-level task-specific scenarios (e.g., recognition) rather than finetuning generative models. Additionally, we devise a suitable regularization strategy to retain the strong generalizability that elucidates the training efficiency and generalizability preservation of our method.
@@
Extensive experiments demonstrate the effectiveness of our OrthCR by evaluating on representative benchmarks: base-to-base/base-to-new, cross-dataset transfer and domain generalization. In the base-to-base/base-to-new setting, our method improves the new class of baseline model by 13.3% on average across 11 datasets, by 0.95% for cross-dataset setting and 1.80% on average across the four datasets for domain generalization setting, all of which presents competitive performance over the existing SoTAs. In summary, our contributions can be summarized as follows:
@@
  • We introduce a novel and efficient orthogonal finetuning method to adapt the VLMs into task-specific knowledge while maintaining strong generalizability. Due to the norm-preserving property, this finetuning leads to stable and faster convergence and exhibits superiority over the prompt tuning methods.
@@
  • To further mitigate the deviation from the pretrained model, we design a Cross-Regularization strategy to enforce the finetuned model distilling informative zero-shot generalization information of the pretrained logits.
@@
  • Cutout data augmentation is employed to enhance the task-specific knowledge when finetuning the VLM under data-efficient setting.
  • Extensive experiments are conducted to validate the effectiveness and effciency of our method, including base-to-base/base-to-new, cross-dataset transfer and domain generalization evaluations.
@@ 2 Related works
Vision language models. Recently, with a significant upsurge of large-scale pretrained vision-language models (VLMs) [81; 86; 35; 12; 64; 72], text and image embeddings have been trained jointly to be aligned with the large-scale image-text pairs corpora. Driven by contrastive loss in a self-supervised manner, VLMs like CLIP [64], ALIGN [35], LiT [84], FLIP [46] and Florence [81] have elucidated remarkable performance. For instance, CLIP [64] and ALIGN [35] utilize approximately 400 million and 1 billion image-text pairs, respectively, to accomplish their multi-modal alignment training, benefiting a wide spectrum of downstream vision and vision-language tasks, including few-shot image-level recognition [87; 88; 85; 21; 37; 89; 68; 56; 11; 74], object detection [20; 24; 3; 82] and segmentation [17; 6; 65; 76]. Despite strong generalizability towards zero-shot recognition tasks of these VLMs, effectively transferring them to downstream tasks without degrading their inherent generalization ability remains a challenging problem.
@@
Efficient tuning for vision language models. With the emergence of VLMs, efficiently adapting these models to specific downstream tasks with limited data samples has garnered significant interest. Prompt Tuning is firstly proposed in the NLP field [48; 22; 45; 41], which attempts to learn task-specific prompt templates. Recently, in the computer vision community, CoOp [88] pioneers the study by tuning the contextual tokens in text branch of CLIP into a set of learnable tokens to few-shot image recognition, which is further improved by CoCoOp [87] through a Meta-Network [57] paradigm to address the overfitting issue on base classes while generalizing better on unseen classes. To efficiently adapt large pretrained Vision Transformers, VPT [36] and Visual Prompting [2] both insert trainable tokens into the input space of transformer model. To leverage additional prompt learning for dual-branch models like CLIP, a plethora of works  [37; 38; 13; 83; 59; 89; 54; 74] have been proposed to learn these prompts towards a way that treats them as continuous learnable vectors while keeping the original model parameters frozen to retain the strong generalizability. Very recently, Test-Time Prompting [69; 68] emerges with the objective of enforcing consistency regularization between multiply views of a test sample by minimizing their averaged entropy. Another line of work [8; 16; 25] focuses on tuning VLMs over the pretrained weights. Adaptation methods [31; 32; 61] have become increasingly ubiquitous. The LoRA series [32; 49; 15] is widely used to finetune pretrained model weights using low-rank matrix optimization. Our method shares a similar principle with LoRA for adapting pretrained model weights, but introduces a novel Orthogonality Learning approach. This not only enhances performance for specific downstream tasks (e.g., few-shot recognition) but also improves robustness and generalization with more efficient convergence.
Orthogonality regularization. Orthogonality has been commonly adopted to introduce orthogonal regularization to improve the robustness of Deep Neural Networks [53; 7; 33; 80; 34; 42; 1; 77; 62; 44], that norm-preserving property can avoid exploding or vanishing gradients during training [4; 23], leading to faster convergence and encouraging robustness and generalization. This objective can be reached by a simple Cayley parameterization [28; 42]. Recently,OPT [51] introduces an orthogonal transformation applied to the neural weights to maintain the minimum hyperspherical energy. Furthermore, OFT [63] extend this orthogonal paradigm to finetune the text-to-image diffusion models by employing Cayley parameterization constraint during the finetuning. In this paper, we further explore the utilization of orthogonal finetuning on CLIP for specific downstream tasks while proposing different regularization strategies to enhance generalizability on novel/uneen classes.
@@ 3.1 Preliminaries
Contrastive Language-Image Pre-training (CLIP). CLIP consists of two parallel encoders, image and text encoders, represented by $\mathcal{\theta}_{CLIP}=\{{\theta_{v},\theta_{t}}\}$. The image encoder $\mathcal{F}_{v}$ can be either a CNN [26] or a ViT [73; 18] for mapping input image into a image embedding, and the text encoder $\mathcal{F}_{t}$ is a Transformer [16] for mapping input text into a text embedding, respectively. During pre-training, CLIP utilizes two parallel encoders to separately encode image and text into corresponding vectors in jointly aligned embedding space, and then adopts contrastive loss to pull together the cosine similarities of the correct image-text vector pairs while pushing away the cosine similarities of incorrect pairs. After pretrained on large-scale image-text pairs corpora, CLIP is capable of computing the text-image similarity and can be generalized to downstream tasks, like zero-shot image recognition, without fine-tuning. Specifically, the input image ${X}$ is first divided into $M$ patches and then projected into patch tokens, and a global class token $[CLS]$ is prepended to the patch token sequence, obtaining ${X}^{~{}}_{0}=\{CLS,e_{1},e_{2},...,e_{M}\}$ where $e_{i}$ standds for the ${i}^{th}$ patch. Those patch tokens will be encoded by transformer blocks inside the image encoder $\mathcal{F}_{v}$ by ${f}^{~{}}_{v}={\mathcal{F}}_{v}({X}^{~{}}_{0}:\theta_{v})$. Given the labels $\{[class]_{c}\}^{C}_{c=1}$ for the $C$ categories for classification where $[class]_{c}$ represents the class name of the $c^{th}$ class, a hand-crafted text prompt like ‘a photo of a $[CLS]$’ will be embedded within the class label $[class]_{c}$ This results in $\mathcal{Y}^{~{}}_{0}=\{SOS,t_{1},t_{2},...,t_{L},c_{k},EOS\}$ where $SOS$ and $EOS$ denote the start and end token embeddings while $t_{i}$ and $c_{k}$ are specific word embedding corresponding to the text prompt and the class label, respectively. The text encoder $\mathcal{F}_{t}$ will encode $\mathcal{Y}^{~{}}_{0}$ via transformer blocks to produce text feature embeddings as ${f}^{~{}}_{t}=\mathcal{F}_{t}(\mathcal{Y}^{~{}}_{0}:\theta_{t})$. During zero-shot inference, the prediction probability on image ${X}$ will be computed as ${p}(y_{i}|X)=\frac{exp(sim(f_{t}\cdot f_{v})/\tau)}{\sum^{C}_{i=1}exp(sim(f^{~{}}_{t}\cdot f^{~{}}_{v})/\tau)}$, where $\tau$ is a learned temperature coefficient and $sim$ denotes the cosine similarity computation, respectively.
@@
Context Optimization (CoOp) [88] proposes to leverage tunable text prompt by replacing the cumbersome and fixed hand-crafted prompt, that can be learnt from data. Now, the tunable prompt is constructed with $M$ learnable continues context vectors as $w=\{w_{1},w_{2},...,w_{M},c_{k}\}$, where $w_{i}$ represents the $i^{th}$ tunable vector and $c_{k}$ denotes the $c^{th}$ class name $[class]_{c}$. The finally finetuned training objective of CoOp is to optimize the contextual vectors $w_{i}$ only by minimize the cross-entropy loss between the ground-truth $\hat{y}$ and the model prediction $y$ as:
@@
3.2 Orthogonal finetuning
Traditionally, finetuning VLMs into specific downstream scenarios typically embraces small learning rate with gradient descent optimizer to update the model, This scheme implicitly constrains risky deviation from pretrained model, aiming to finetune the model via implicitly minimizing $\|\bm{M}-\bm{M}_{0}\|$ where $\bm{M}$ is the finetuned model weights and $\bm{M}_{0}$ is the pretrained model weights. Towards this strategy, there are still various ways to finetune a pretrained VLM. For example, LoRA [32] employs an additive low-rank matrix with constraint for model weights update, i.e., $\text{rank}(\bm{M}-\bm{M}_{0})=r^{\prime}$ where $r^{\prime}$ is set to be relatively smaller number than the pretrained ones. Differently, Orthogonal transformation targets at inducing a constraint for the pairwise similarity between neurons [51; 63]: $\|\text{HE}(\bm{M})-\text{HE}(\bm{M}_{0})\|=0$ ,where $\text{HE}(\cdot)$ denotes hyperspherical energy of a weight matrix. In this paper, we draw attention to the Feed-Forward-Networks (FFN) within the transformer architecture of CLIP, shown in Fig 2. Suppose a fully-connected layer with $\bm{W}=\{\bm{w}_{1},\cdots,\bm{w}_{n}\}\in\mathbb{R}^{d\times n}$ where $\bm{w}_{i}\in\mathbb{R}^{d}$ is the $i^{th}$ neuron ($\bm{W}_{0}$ is the pretrained weights). We expect to acquire the output vector $\bm{z}\in\mathbb{R}^{n}$ by $\bm{z}=\bm{W}^{\top}\bm{x}$ where $\bm{x}\in\mathbb{R}^{d}$ is the input vector. When introducing the orthogonal finetuning as minimizing the hysperical energy difference between the finetuned and pretrained model:

### v3

# Enhancing Robustness of Vision-Language Models through Orthogonality Learning and Self-Regularization

@@ Abstract
Efficient fine-tuning of vision-language models (VLMs) like CLIP for specific downstream tasks is gaining significant attention. Previous works primarily focus on prompt learning to adapt the CLIP into a variety of downstream tasks, however, suffering from task overfitting when fine-tuned on a small data set. In this paper, we introduce an orthogonal fine-tuning method for efficiently fine-tuning pretrained weights and enabling enhanced robustness and generalization, while a self-regularization strategy is further exploited to maintain the stability in terms of zero-shot generalization of VLMs, dubbed OrthSR. Specifically, trainable orthogonal matrices are injected seamlessly into the transformer architecture and enforced with orthogonality constraint during the training, benefiting from the norm-preserving property and thus leading to stable and faster convergence, while keeping the pre-trained weights frozen. To alleviate deviation from fine-tuning, a self-regularization strategy is further employed to retain the generalization of the model during the training within a bypass manner. In addition, to enrich the sample diversity for downstream tasks under the small dataset scenario, we first explore attentive CutOut data augmentation to boost the efficient fine-tuning, leading to better model fitting capacity for specific downstream task. Then we support the theoretical analysis on how our approach improves the specific downstream performance and maintains the generalizability. For the first time, we revisit the CLIP and CoOp with our method to effectively improve the model on few-shot image classficiation scenario on par with the elaborated prompt learning methods. We conduct extensive experiments to demonstrate that our method explicitly steers pretrained weight space to represent the task-specific knowledge and presents competitive generalizability under base-to-base/base-to-new, cross-dataset transfer and domain generalization evaluations.
@@ 1 Introduction
Large-scale pre-trained vision-language models (VLMs) have been emerging as prevalent cornerstones in a wide spectrum of downstream vision and vision-language tasks, including few-shot image recognition Zhou et al. (2022a; b); Zhang et al. (2022a); Gao et al. (2024); Khattak et al. (2023a); Zhu et al. (2023); Samadh et al. (2023); Manli et al. (2022); Cheng et al. (2023); Wang et al. (2023), object-detection Feng et al. (2022); Gu et al. (2021); Bangalath et al. (2022); Zang et al. (2022a) and segmentation Ding et al. (2022); Boyi et al. (2022); Rao et al. (2022); Wang et al. (2024). Leading models like CLIP Radford et al. (2021) and ALIGN Jia et al. (2021) demonstrate remarkable generalizability by training with aligning image-text pairs from large web corpora using contrastive loss, thereby encoding open-vocabulary concepts within a joint vision-language embedding space. Despite the effectiveness of these VLMs in zero-shot recognition, fine-tuning them for specific downstream tasks while preserving their strong zero-shot capabilities remains a significant challenge. Designing manual text prompts for different tasks requires substantial human effort and expert knowledge, which is often infeasible for achieving optimal performance in data-efficient settings Brown et al. (2020).
@@
Recently, prompt learning Zhou et al. (2022b; a) serves as an exceptional paradigm to achieve this objective, however, tending to prioritize task-specific knowledge and resulting in task overfitting issues Park et al. (2024); Khattak et al. (2023b), where the fine-tuned model struggles to generalize well to new/unseen tasks under data-efficient settings. To address this dilemma, alternative approaches must be explored. Drawing inspiration from empirical observations that hyperspherical similarity effectively encodes semantic information Chen et al. (2020); Liu et al. (2018b; 2017) and that hyperspherical energy Liu et al. (2018a) can characterize the pairwise relational structure among neurons, we hypothesize that well-pretrained models like CLIP should maintain consistent levels of hyperspherical energy even after fine-tuning. An intuitive approach is to use a suitable regularizer to preserve hyperspherical energy levels during the fine-tuning phase. However, ensuring that the difference in hyperspherical energy is minimized remains a challenge. Inspired by recent orthogonal transformation methods Qiu et al. (2023); Liu et al. (2021), we propose that the pretrained pairwise hyperspherical energy can be preserved by leveraging orthogonal transformation for all neurons with the same operation. This approach utilizes the invariance property of orthogonal transformation, meaning norm-preserving during fine-tuning, to maintain consistent hyperspherical energy levels.
@@
Motivated by the preservation of hyperspherical energy through orthogonal transformation, we introduce Orthogonality Learning to adapt pretrained VLMs (e.g., CLIP) to specific downstream tasks (e.g., few-shot image recognition) without altering their hyperspherical energy, thanks to the norm-preserving property during fine-tuning. This approach differs from common methods that heavily rely on prompt learning. Furthermore, previous works Lin et al. (2020); Liu et al. (2018a; 2021) have shown that small hyperspherical energy leads to better generalization, and orthogonal transformation is a suitable and flexible solution for achieving this, especially in classification task. Our main idea is to apply the same orthogonal transformation to neurons so that pairwise angles are maintained within the hypersphere of CLIP. Although prevalent adaptation methods for pretrained weights, such as LoRA Hu et al. (2022), achieve fine-tuning by adding small component matrices, they still suffer from low training convergence and generalizability degradation.
Figure 1: The pipeline comparison for tuning or adapting VLMs into downstream tasks. Our contribution is to introduce a new fine-tuning pipeline by orthogonal tuning, that boost the CLIP and CoOp with competitive base/novel accuracy performances when compared with existing methods (results are computed by average 11 datasets).
@@
In this paper, we propose a novel and efficient fine-tuning method using Orthogonality Learning, motivated by the preservation of hyperspherical energy through orthogonal transformation, shown different paradigm with exisitng works in Fig. 1 (a). To mitigate deviation from orthogonal constraint during training, we introduce a Self-Regularization strategy using the initial pretrained weights as an anchor point, thus dubbed OrthSR. Our method keeps the pretrained weights frozen while applying orthogonal fine-tuning and regularization simultaneously. In the dual-branch transformer architecture of the CLIP model, we inject trainable orthogonal matrices and enforce orthogonal constraints (such as using Cayley parameterization Helfrich et al. (2018); Lezcano-Casado & Martınez-Rubio (2019)). This ensures each injected layer matrix is orthogonal with a determinant of 1. We investigate orthogonal fine-tuning in both image and text encoder of CLIP to demonstrate training efficiency and generalizability preservation of our method, distinguishing it from prompt tuning and low-rank matrix decomposition methods. The norm-preserving property of orthogonal transformations helps maintain hyperspherical energy levels, benefiting of stable convergence, robustness, and generalization. This enables seamless integration of task-specific knowledge into pretrained VLMs, allowing the trainable matrices to be merged with frozen weights during deployment without adding inference latency, while we shows evaluation superiority over previous methods in Fig. 1 (b). To prevent significant deviations from the pretrained model, we employ a Self-Regularization strategy that guides the model to stay close to the anchor point, supported by the pretrained model within a bypass manner. This simple yet effective approach sustains orthogonal fine-tuning with initial anchor regularization, avoiding deviations from the zero-shot generalizability manifold severely. Besides, we utilize attentive CutOut data augmentation to enrich the data diversity, enhancing the task-specific knowledge of fine-tuned model (e.g., few-shot image recognition) under data-efficient setting. This leads to better model fitting capacity for specific downstream task, serving as implicitly increasing the sample diversity. Unlike previous works Qiu et al. (2023); Liu et al. (2021), we focus on adapting VLMs to high-level task-specific scenarios (e.g., recognition) rather than fine-tuning generative models. Additionally, we devise a suitable regularization strategy to retain the strong generalizability that elucidates the training efficiency and generalizability preservation of our method.
@@
Extensive experiments demonstrate the effectiveness of our OrthSR by evaluating on representative benchmarks: base-to-base/base-to-new, cross-dataset transfer and domain generalization. In the base-to-base/base-to-new setting, our method improves the new class of baseline model by 13.3% on average across 11 datasets, by 0.95% for cross-dataset setting and 1.80% on average across the four datasets for domain generalization setting, all of which presents competitive performance over the existing SoTAs. In summary, our contributions can be summarized as follows:
@@
  • We introduce a novel and efficient orthogonal fine-tuning method to adapt the VLMs into task-specific knowledge while maintaining strong generalizability. Due to the norm-preserving property, this fine-tuning leads to stable and faster convergence and exhibits superiority over the prompt tuning methods.
@@
  • To further mitigate the deviation from the pretrained model, we design a Self-Regularization strategy to enforce the fine-tuned model distilling informative zero-shot generalization information of the pretrained logits.
@@
  • Attentive CutOut data augmentation is employed to enhance the task-specific knowledge when fine-tuning the VLM under data-efficient setting.
  • Extensive experiments are conducted to validate the effectiveness and effciency of our method, for the first time, we boost the CLIP and CoOp with weight decomposition tuning to obtain on par or even superior performances over existing methods.
@@ 2 Related works
Vision language models. Recently, with a significant upsurge of large-scale pretrained vision-language models (VLMs) Yuan et al. (2021); Zhang et al. (2022b); Jia et al. (2021); Cherti et al. (2023); Radford et al. (2021); Touvron et al. (2023), text and image embeddings have been trained jointly to be aligned with the large-scale image-text pairs corpora. Driven by contrastive loss in a self-supervised manner, VLMs like CLIP Radford et al. (2021), ALIGN Jia et al. (2021), LiT Zhai et al. (2022), FLIP Li et al. (2023b) and Florence Yuan et al. (2021) have elucidated remarkable performance. For instance, CLIP Radford et al. (2021) and ALIGN Jia et al. (2021) utilize approximately 400 million and 1 billion image-text pairs, respectively, to accomplish their multi-modal alignment training, benefiting a wide spectrum of downstream vision and vision-language tasks, including few-shot image-level recognition Zhou et al. (2022a; b); Zhang et al. (2022a); Gao et al. (2024); Khattak et al. (2023a); Zhu et al. (2023); Samadh et al. (2023); Manli et al. (2022); Cheng et al. (2023); Wang et al. (2023), object detection Feng et al. (2022); Gu et al. (2021); Bangalath et al. (2022); Zang et al. (2022a) and segmentation Ding et al. (2022); Boyi et al. (2022); Rao et al. (2022); Wang et al. (2024). Despite strong generalizability towards zero-shot recognition tasks of these VLMs, effectively transferring them to downstream tasks without degrading their inherent generalization ability remains a challenging problem.
@@
Efficient tuning for vision language models. With the emergence of VLMs, efficiently adapting these models to specific downstream tasks with limited data samples has garnered significant interest. Prompt Tuning is firstly proposed in the NLP field Liu et al. (2023); Gao et al. (2021); Li & Liang (2021); Lester et al. (2021), which attempts to learn task-specific prompt templates. Recently, in the computer vision community, CoOp Zhou et al. (2022b) pioneers the study by tuning the contextual tokens in text branch of CLIP into a set of learnable tokens to few-shot image recognition, which is further improved by CoCoOp Zhou et al. (2022a) through a Meta-Network Munkhdalai & Yu (2017) paradigm to address the overfitting issue on base classes while generalizing better on unseen classes. To efficiently adapt large pretrained Vision Transformers, VPT Jia et al. (2022) and Visual Prompting Bahng et al. (2022) both insert trainable tokens into the input space of transformer model. To leverage additional prompt learning for dual-branch models like CLIP, a plethora of works  Khattak et al. (2023a; b); Cho et al. (2023); Zang et al. (2022b); Park et al. (2024); Zhu et al. (2023); Lu et al. (2022); Wang et al. (2023) have been proposed to learn these prompts towards a way that treats them as continuous learnable vectors while keeping the original model parameters frozen to retain the strong generalizability. Very recently, Test-Time Prompting Shu et al. (2022); Samadh et al. (2023) emerges with the objective of enforcing consistency regularization between multiply views of a test sample by minimizing their averaged entropy. Another line of work Brown et al. (2020); Devlin et al. (2019); He et al. (2022) focuses on tuning VLMs over the pretrained weights. Adaptation methods Houlsby et al. (2019); Hu et al. (2022); Poth et al. (2023) have become increasingly ubiquitous. The LoRA series Hu et al. (2022); Liu et al. (2024); Dettmers et al. (2024) is widely used to finetune pretrained model weights using low-rank matrix optimization. Our method shares a similar principle with LoRA for adapting pretrained model weights, but introduces a novel Orthogonality Learning approach. This not only enhances performance for specific downstream tasks (e.g., few-shot recognition) but also improves robustness and generalization with more efficient convergence. Orthogonality regularization. Orthogonality has been commonly adopted to introduce orthogonal regularization to improve the robustness of Deep Neural Networks Liu et al. (2017); Brock et al. (2017); Huang et al. (2020); Xie et al. (2017); Huang et al. (2018); Lezcano-Casado & Martınez-Rubio (2019); Arjovsky et al. (2016); Wisdom et al. (2016); Qi et al. (2020); Li et al. (2019), that norm-preserving property can avoid exploding or vanishing gradients during training Bengio et al. (1994); Glorot & Bengio (2010), leading to faster convergence and encouraging robustness and generalization. This objective can be reached by a simple Cayley parameterization Helfrich et al. (2018); Lezcano-Casado & Martınez-Rubio (2019). Recently,OPT Liu et al. (2021) introduces an orthogonal transformation applied to the neural weights to maintain the minimum hyperspherical energy. Furthermore, OFT Qiu et al. (2023) extend this orthogonal paradigm to finetune the text-to-image diffusion models by employing Cayley parameterization constraint during the finetuning. In this paper, we further explore the utilization of orthogonal finetuning on CLIP for specific downstream tasks while proposing different regularization strategies to enhance generalizability on novel/uneen classes.
@@ 3.1 Preliminaries
Contrastive Language-Image Pre-training (CLIP). CLIP consists of two parallel encoders, image and text encoders, represented by $\mathcal{\theta}_{CLIP}=\{{\theta_{v},\theta_{t}}\}$. The image encoder $\mathcal{F}_{v}$ can be either a CNN He et al. (2016) or a ViT Vaswani et al. (2017); Dosovitskiy et al. (2020) for mapping input image into a image embedding, and the text encoder $\mathcal{F}_{t}$ is a Transformer Devlin et al. (2019) for mapping input text into a text embedding, respectively. During pre-training, CLIP utilizes two parallel encoders to separately encode image and text into corresponding vectors in jointly aligned embedding space, and then adopts contrastive loss to pull together the cosine similarities of the correct image-text vector pairs while pushing away the cosine similarities of incorrect pairs. After pretrained on large-scale image-text pairs corpora, CLIP is capable of computing the text-image similarity and can be generalized to downstream tasks, like zero-shot image recognition, without fine-tuning. Specifically, the input image ${X}$ is first divided into $M$ patches and then projected into patch tokens, and a global class token $[CLS]$ is prepended to the patch token sequence, obtaining ${X}^{~{}}_{0}=\{CLS,e_{1},e_{2},...,e_{M}\}$ where $e_{i}$ standds for the ${i}^{th}$ patch. Those patch tokens will be encoded by transformer blocks inside the image encoder $\mathcal{F}_{v}$ by ${f}^{~{}}_{v}={\mathcal{F}}_{v}({X}^{~{}}_{0}:\theta_{v})$. Given the labels $\{[class]_{c}\}^{C}_{c=1}$ for the $C$ categories for classification where $[class]_{c}$ represents the class name of the $c^{th}$ class, a hand-crafted text prompt like ‘a photo of a $[CLS]$’ will be embedded within the class label $[class]_{c}$ This results in $\mathcal{Y}^{~{}}_{0}=\{SOS,t_{1},t_{2},...,t_{L},c_{k},EOS\}$ where $SOS$ and $EOS$ denote the start and end token embeddings while $t_{i}$ and $c_{k}$ are specific word embedding corresponding to the text prompt and the class label, respectively. The text encoder $\mathcal{F}_{t}$ will encode $\mathcal{Y}^{~{}}_{0}$ via transformer blocks to produce text feature embeddings as ${f}^{~{}}_{t}=\mathcal{F}_{t}(\mathcal{Y}^{~{}}_{0}:\theta_{t})$. During zero-shot inference, the prediction probability on image ${X}$ will be computed as ${p}(y_{i}|X)=\frac{exp(sim(f_{t}\cdot f_{v})/\tau)}{\sum^{C}_{i=1}exp(sim(f^{~{}}_{t}\cdot f^{~{}}_{v})/\tau)}$, where $\tau$ is a learned temperature coefficient and $sim$ denotes the cosine similarity computation, respectively.
@@
Context Optimization (CoOp) Zhou et al. (2022b) proposes to leverage tunable text prompt by replacing the cumbersome and fixed hand-crafted prompt, that can be learnt from data. Now, the tunable prompt is constructed with $M$ learnable continues context vectors as $w=\{w_{1},w_{2},...,w_{M},c_{k}\}$, where $w_{i}$ represents the $i^{th}$ tunable vector and $c_{k}$ denotes the $c^{th}$ class name $[class]_{c}$. The finally fine-tuned training objective of CoOp is to optimize the contextual vectors $w_{i}$ only by minimize the cross-entropy loss between the ground-truth $\hat{y}$ and the model prediction $y$ as:
@@
3.2 Orthogonal fine-tuning
Traditionally, fine-tuning VLMs into specific downstream scenarios typically embraces small learning rate with gradient descent optimizer to update the model, This scheme implicitly constrains risky deviation from pretrained model, aiming to finetune the model via implicitly minimizing $\|\bm{M}-\bm{M}_{0}\|$ where $\bm{M}$ is the fine-tuned model weights and $\bm{M}_{0}$ is the pretrained model weights. Towards this strategy, there are still various ways to finetune a pretrained VLM. For example, LoRA Hu et al. (2022) employs an additive low-rank matrix with constraint for model weights update, i.e., $\text{rank}(\bm{M}-\bm{M}_{0})=r^{\prime}$ where $r^{\prime}$ is set to be relatively smaller number than the pretrained ones. Differently, Orthogonal transformation targets at inducing a constraint for the pairwise similarity between neurons Liu et al. (2021); Qiu et al. (2023): $\|\text{HE}(\bm{M})-\text{HE}(\bm{M}_{0})\|=0$ ,where $\text{HE}(\cdot)$ denotes hyperspherical energy of a weight matrix. In this paper, we draw attention to the Feed-Forward-Networks (FFN) within the transformer architecture of CLIP, shown in Fig 2. Suppose a fully-connected layer with $\bm{W}=\{\bm{w}_{1},\cdots,\bm{w}_{n}\}\in\mathbb{R}^{d\times n}$ where $\bm{w}_{i}\in\mathbb{R}^{d}$ is the $i^{th}$ neuron ($\bm{W}_{0}$ is the pretrained weights). We expect to acquire the output vector $\bm{z}\in\mathbb{R}^{n}$ by $\bm{z}=\bm{W}^{\top}\bm{x}$ where $\bm{x}\in\mathbb{R}^{d}$ is the input vector. When introducing the orthogonal fine-tuning as minimizing the hysperical energy difference between the fine-tuned and pretrained model: